OpenStudy (anonymous):

What is the slope of the line 2x + 1/3y = –8? A. –6 B. –24 C. 2 D. -1/3

ganeshie8 (ganeshie8):

write it in slope intercept form

OpenStudy (anonymous):

1/3y = -2x -8

OpenStudy (anonymous):

is it correct?

ganeshie8 (ganeshie8):

slope intercept form is y = mx + c

ganeshie8 (ganeshie8):

just multiply whole equation with 3

OpenStudy (anonymous):

y= -2/3x - 8/3

OpenStudy (anonymous):

is it correct?

ganeshie8 (ganeshie8):

1/3y = -2x -8 multiply 3 3(1/3y) = 3(-2x -8 )

ganeshie8 (ganeshie8):

y = -6x -24

ganeshie8 (ganeshie8):

So... slope = ?

OpenStudy (anonymous):

-6?

ganeshie8 (ganeshie8):

Yes !

OpenStudy (anonymous):

Thank you!!!

ganeshie8 (ganeshie8):

y = mx + c ^ when you want slope, just write it like above. and x-coefficent is the slope

ganeshie8 (ganeshie8):

yw :)

OpenStudy (anonymous):

THank you!